Find five consecutive odd integers such that the sum of the first and the fifth is one less than three times the fourth.
Answer by checkley71(8403)

X
X+2
X+4
X+6
X+8
X+(X+8)=3(X+6)-1
2X+8=3X+18-1
2X-3X=17-8
-X=9
X=-9
X+2=-7
X+4=-5
X+6=-3
X+8=-1
PROOF
-9+(-9+8)=3(-9+6)-1
-9-1=3*-3-1
-10=-10